OpenStudy (anonymous):

write an equation of the line that passes through (1,-6) and is perpendicular to the line y=3x+7

OpenStudy (abb0t):

The relationship for the slopes of perpendicular lines is: \(m_1 \times m_2 = -1\) You already have \(m_1\) but you don't have \(m_2\) so solve for it. If you have \(3m_2=-1\), what is \(m_2\)? Now, you can use the point-slope equation: \(y - y_1 = m(x - x_1)\) which you should be familiar with. To get the new equation that's perpendicular to your first line.

OpenStudy (abb0t):

remember, perpendicular means \(\perp\)

OpenStudy (abb0t):

You could also use calculus to solve this. Simply find the derivative of your function y = mx+ b with respect to "x". Then, use the inverse of the slope (to get your perpendicular). Then follow the same steps, plug in your new slope into your point-slope equation solvinf rot the new equation. Good Luck. Cheers